calculate the internal energy of enthalpy of any random system.also find out the mathematical relation between internal energy enthalpy and then entropy?


1
Expert's answer
2021-08-03T05:14:42-0400



"U=\frac{m}{M}C_VT"


For air "M=28.96 \cdot 10^{-3}kg/mol" and "C_V=\frac{i}{2}R=\frac{5}{2}\cdot 8.31\approx20.8\frac {J}{mol\cdot K}"


"U=\frac{m}{M}C_VT=\frac{1}{28.96 \cdot 10^{-3}}\cdot 20.8\cdot 293=210442J"


The enthalpy of air


"H=U+pV=210442+2\cdot 10^6\cdot 0.05=310442J"


Enthalpy is the sum total of all the energies, whereas entropy is the measure of the change in enthalpy/temperature.